### v1

@@ Abstract
Political scientists are increasingly attuned to the promises and pitfalls of establishing causal effects. But the vital question for many is not if a causal effect exists but why and how it exists. Even so, many researchers avoid causal mediation analyses due to the assumptions required, instead opting to explore causal mechanisms through what we call intermediate outcome tests. These tests use the same research design used to estimate the effect of treatment on the outcome to estimate the effect of the treatment on one or more mediators, with authors often concluding that evidence of the latter is evidence of a causal mechanism. We show in this paper that, without further assumptions, this can neither establish nor rule out the existence of a causal mechanism. Instead, such conclusions about the indirect effect of treatment rely on implicit and usually very strong assumptions that are often unmet. Thus, such causal mechanism tests, though very common in political science, should not be viewed as a free lunch but rather should be used judiciously, and researchers should explicitly state and defend the requisite assumptions.
@@ 1 Introduction
There has long been disagreement about the proper way to test for causal mechanisms, leading to a fractured landscape and difficulty in understanding how different tests relate. From a formal point of view, causal mechanisms have been most closely linked to mediation analysis, wherein the overall average treatment effect is decomposed into its direct (net of a mediating variable of interest) and indirect effects (Pearl, 2001; Imai, Keele and Yamamoto, 2010; Imai et al., 2011). Some scholars, however, object to causal mediation analysis because of the strong assumptions it requires (Green, Ha and Bullock, 2010; Acharya, Blackwell and Sen, 2016; Bullock and Green, 2021; Callis, Dunning and Tuñón, 2024). In particular, identification of the average direct and indirect effects relies on unverifiable assumptions that (i) the mediator is as-if randomized condition on pretreatment covariates, and (ii) there are no other post-treatment variables that confound the mediator-outcome relationship. This skepticism is reflected in the relative lack of causal mediation analyses in recent empirical political science articles. We analyzed 487 empirical papers published in the American Political Science Review, the American Journal of Political Science, and The Journal of Politics between 2022 and 2023 and found that 161 (33%) provided at least one quantitative test for causal mechanisms. As shown by Figure 1, however, of these 161 papers, only 16 (10%) employed a formal causal mediation analysis.
If these studies do not employ formal techniques of causal mediation, how do they assess causal mechanisms? Figure 1 shows it is common for researchers to employ an alternative approach to evaluating causal mechanisms, which in this paper we call intermediate outcome tests or IOTs. In these tests, researchers estimate the causal effect of the treatment on one or more potential mediators, often using the same research design used to estimate the main effect of the treatment on the outcome (Green, Ha and Bullock, 2010; Callis, Dunning and Tuñón, 2024). Under this approach, if researchers detect an average treatment effect on a mediator, this mediator is considered part of a potential “mechanism” for the treatment effect on outcome. IOTs have clear, intuitive appeal since it seems to follow that, for $M$ to mediate the relationship between treatment $A$ and outcome $Y$, there must be an effect of $A$ on $M$ and then of $M$ on $Y$. The (often implicit) argument for the IOT approach is that an average effect of $A$ on $M$ is a necessary and sometimes sufficient condition for $M$ to be a causal mechanism. Thus, it is no surprise that this simple ad hoc approach is one of the dominant ways to estimate causal mechanisms, accounting for roughly 40% of the papers in our literature review.
@@
In this paper, we show that, in an experimental context, intermediate outcome tests under randomization of treatment alone can neither establish nor rule out the existence of an indirect effect of treatment on the outcome through that mediator. Papers that imply otherwise are engaging (often unknowingly) in what we call assumption smuggling or presenting suspect conclusions that rely on strong hidden assumptions alongside well-grounded findings that depend on the more reasonable, explicitly stated assumptions. We establish the IOT’s inability to inform mechanisms using sharp nonparametric bounds that contain all values of the indirect effect consistent with the observed data and the maintained assumptions. Under randomization of treatment alone, these bounds are never informative about the sign of the indirect effect for two reasons. First, the average effect of the treatment on the mediator might mask heterogeneous treatment effects that are important to the causal mechanism. Second, establishing indirect effects requires strong evidence or assumptions about the effect of the mediator on the outcome at the individual level.
@@
To help researchers navigate this, we further introduce a set of assumptions that, though very rarely discussed in practice, do enable the IOT procedure to have a meaningful interpretation. Specifically, we investigate a monotonic mediator response assumption that the units’ value of the mediator can only be affected by treatment in one direction. Under this assumption, we show that a lack of treatment effect on the mediator can, in fact, rule out indirect effects through that mediator. We also derive the nonparametric sharp bounds for the average indirect effect under this assumption, showing that IOTs can “approximately” rule out mechanisms in that small (zero) effects of the treatment on the outcome imply small (zero) indirect effects. Of course, monotonicity is a strong, untestable assumption that will frequently be violated in applied practice, and there is no design-based way to ensure it holds. Thus, it is critical for researchers to state this assumption clearly and to interrogate it if they wish to use IOTs to assess causal mechanisms at all.
@@
Can IOTs ever establish the sign of an indirect effect? We also show that, unfortunately, it is nearly impossible to establish the sign of an indirect effect even under very strong assumptions. In particular, we show that even if we assume a positive average effect of the mediator on the outcome, the sharp nonparametric bounds on the indirect effect will never be informative about the sign of that effect. Thus, while the monotonic mediator response assumption may allow researchers to rule out potential pathways, no intermediate outcome test can establish the simple existence of a causal mechanism even if we make heroic assumptions about the mediator-outcome relationship.
@@
The contributions of this paper are twofold. First, we highlight for applied researchers that this ad hoc, widely used approach to analyze causal mechanisms actually relies on strong and unusually unstated assumptions; relaxing the assumptions, as we show, can render substantive conclusions invalid. Second, by clarifying the assumptions needed, we provide guidance for applied researchers looking to provide substantive meaning for the interpretation of these IOTs. The takeaway message of this paper reinforces what has been argued elsewhere: establishing causal mechanisms is difficult and requires strong assumptions regardless of the approach. Researchers may not avoid this difficulty even when using IOTs.
@@
Our work contributes to a broader literature on causal mechanisms and (in)direct effects in multiple fields. Many scholars have formalized the concept of direct and indirect effects in terms of potential outcomes and have derived assumptions that can point identify these quantities (Pearl, 2001; Imai, Keele and Yamamoto, 2010). Most similar to our own work is Imai, Tingley and Yamamoto (2013), which derived nonparametric bounds for indirect effects under several different experimental designs, and Glynn (2012), which showed how individual-level heterogeneity in treatment effects renders many indirect effect tests ineffective. We expand on their results by showing how different causal assumptions can change the bounds for a single experiment. Several studies have proposed alternative ways to assess causal mechanisms, such as implicit mediation (Bullock and Green, 2021) and effect heterogeneity (Fu and Slough, 2024). However, these papers show that these alternatives also rely on strong assumptions. An alternative definition of causal mechanisms centers on a quantity known as the “controlled direct effect” (CDE), which captures the average direct effect of a treatment variable while holding the mediator fixed at a given value (Goetgeluk, Vansteelandt and Goetghebeur, 2008; Acharya, Blackwell and Sen, 2016, 2018; Zhou and Wodtke, 2019; Blackwell and Strezhnev, 2022). While identifying the CDE requires weaker assumptions than mediation, many mediation-skeptical scholars object to these approaches because they require us to identify the effect of the mediator on the outcome. [^1] Some recent work investigates how to handle multiple mediating variables (Imai and Yamamoto, 2013; VanderWeele, Vansteelandt and Robins, 2014; Daniel et al., 2015; Zhou and Yamamoto, 2023). These approaches typically rely on stronger identification assumptions than in the single-mediator case, and multiple mechanism tests have yet to be widely adopted in political science. For these reasons, we focus on the single-mediator setting.
@@
This paper proceeds as follows. In Section 2, we discuss the setting of our two empirical applications, showing how researchers use IOTs in practice. Since the phrase “causal mechanisms” means different things to different scholars, we next lay out our preferred notation and definitions for causal effects and causal mechanisms in Section 3. In Section 4, we formally analyze intermediate outcomes tests using a principal stratification approach and derive sharp nonparametric bounds under different assumptions. Section 5 applies these bounds to two recent studies, showing the limited amount that can be learned about indirect effects with these tests. Section 6 discusses how scholars might consider mechanisms when they cannot meet the assumptions we discuss in this paper. Section 7 concludes and describes avenues for future research.
2 Two Motivating Examples
@@ 2.1 Reducing Outgroup Prejudice
An important literature in political psychology examines intergroup prejudices and which forms of individual-oriented interventions might durably reduce such prejudices. An approach that has garnered empirical support in a growing literature is the use of interpersonal conversations (Galinsky and Moskowitz, 2000; Bruneau and Saxe, 2012; Broockman and Kalla, 2016; Adida, Lo and Platas, 2018; Simonovits, Kezdi and Kardos, 2018; Audette, Horowitz and Michelitch, 2020; Lowe, 2021; Williamson et al., 2021), but an unresolved question in this recent work is which “narrative strategies” are most effective at reducing prejudice. For example, an interpersonal conversation that emphasizes taking the perspective of an outgroup member (“perspective taking”) may yield different outcomes from one that simply describes the negative experiences of the outgroup (“ perspective getting”).
@@ 2.2 Transitional Justice Museums and Support for Democracy
Transitional justice policies are frequently credited with facilitating democratic consolidation, reconciliation, and peace-building in post-conflict societies (De Brito, Enríquez and Aguilar, 2001; Horne, 2014; Nalepa, 2010). Less is known, however, about how such policies operate at the individual level and through what mechanisms. In a field experiment in Chile, Balcells, Palanza and Voytas (2022) study one particular instance of transitional justice policy: museums. Students were randomized into either a control condition or a treatment condition consisting of an hour-long tour of a museum memorializing victims of General Augusto Pinochet’s dictatorship. The authors then fielded a survey about beliefs about democracy and found that treated students “display greater support for democratic institutions, are more likely to reject institutions associated with the repressive period, and are more supportive of restorative transitional justice policies" (Balcells, Palanza and Voytas, 2022, p. 496).
@@ 3 Causal mechanisms: notation and quantities of interest
With this notation, it is possible to define potential outcomes that involve “natural” values of the mediator, such as $Y_{i}(1,M_{i}(0))$, which is the outcome that unit $i$ would take if it was assigned to treatment but had received the mediator value under the control condition. These types of potential outcomes are important to mediation, but they create challenges in estimation and interpretation due to their “cross-world” nature. In particular, a unit cannot simultaneously be subject to treatment and control conditions at the same time, so $Y_{i}(1,M_{i}(0))$ is not observable even in principle.
@@ 3.1 Identifying assumptions
Assumption 1 is justified by randomized experiments but is also the key identifying assumption of papers that invoke selection on observables to justify causal interpretations of estimates. While the latter setting is often thought to be less plausible, a large body of work focuses on so-called “natural” experiments that could make this assumption believable. Under this assumption, we can identify the ATE as the average of covariate-specific effects,
@@
Many authors have expressed skepticism about Assumption 2 since it requires that there are no (measured or unmeasured) post-treatment confounders for the relationship between the mediator and the outcome. This means the effect of $M_{i}$ on $Y_{i}$ must be identified using the exact same design as the overall effect of treatment. [^2] One alternative approach that has garnered some attention in recent years involves estimating a quantity known as the controlled direct effect (CDE), which captures the causal effect of a treatment when the mediator is fixed to a given value. This quantity is attractive because it is identified under a weaker assumption than Assumption 2, one that allows for the existence of observed “intermediate confounders” (confounders that are affected by the treatment and which affect the mediator and outcome). A drawback of this approach is that it does not quantify the extent of mediation through $M_{i}$ and only enables a researcher to rule out the existence of alternative mediators other than $M_{i}$ (see Acharya, Blackwell and Sen (2016)). This condition is often difficult to sustain when the mediator is observed at its natural value after treatment, as is commonly the case. Indeed, while many are willing to believe the identifying assumptions for treatment, assuming the same condition for the mediator is much less plausible.

### v2

@@ Abstract
Political scientists are increasingly interested in assessing causal mechanisms, or determining not just if a causal effect exists but also why it occurs. Even so, many researchers avoid formal causal mediation analyses due to their stringent assumptions, instead opting to explore causal mechanisms through what we call intermediate outcome tests. These tests estimate the effect of the treatment on one or more mediators and view such effects as suggestive evidence of a causal mechanism. In this paper, we use nonparametric bounding analysis to show that, without further assumptions, these tests can neither establish nor rule out the existence of a causal mechanism. To use intermediate outcome tests as a falsification test of causal mechanisms, researchers must make a very strong but rarely discussed monotonicity assumption. We develop a way to assess the plausibility of this monotonicity assumption and estimate our bounds for two recent experiments that use these tests.
@@ 1 Introduction
There has long been disagreement about the proper way to test for causal mechanisms, leading to a fractured landscape and difficulty understanding how different tests relate. From a formal point of view, causal mechanisms have been most closely linked to mediation analysis, wherein the overall average treatment effect is decomposed into its direct (net of a mediating variable of interest) and indirect effects (Pearl, 2001; Imai, Keele and Yamamoto, 2010; Imai et al., 2011). Some scholars, however, object to causal mediation analysis because of the strong assumptions it requires (Green, Ha and Bullock, 2010; Acharya, Blackwell and Sen, 2016; Bullock and Green, 2021; Callis, Dunning and Tuñón, 2024). In particular, identification of the average direct and indirect effects relies on unverifiable assumptions that (i) the mediator is as-if randomized condition on pretreatment covariates, and (ii) there are no other post-treatment variables that confound the mediator-outcome relationship. This skepticism is reflected in the relative lack of causal mediation analyses in recent empirical political science articles. We analyzed 487 empirical papers published in the American Political Science Review, the American Journal of Political Science, and The Journal of Politics between 2022 and 2023 and found that 161 (33%) provided at least one quantitative test for causal mechanisms. As shown by Figure 1, however, of these 161 papers, only 16 (10%) employed a formal causal mediation analysis.
If these studies do not employ formal techniques of causal mediation, how do they assess causal mechanisms? Figure 1 shows that researchers commonly use an alternative approach to evaluating causal mechanisms, which we call in this paper we call intermediate outcome tests or IOTs. In these tests, researchers estimate the causal effect of the treatment on one or more potential mediators, often using the same research design used to estimate the main effect of the treatment on the outcome (Green, Ha and Bullock, 2010; Callis, Dunning and Tuñón, 2024). Under this approach, if researchers detect an average treatment effect on a mediator, this mediator is considered part of a potential “mechanism” for the treatment effect on outcome. IOTs have clear, intuitive appeal since it seems to follow that, for $M$ to mediate the relationship between treatment $A$ and outcome $Y$, there must be an effect of $A$ on $M$ and then of $M$ on $Y$. The (often implicit) argument for the IOT approach is that an average effect of $A$ on $M$ is a necessary and sometimes sufficient condition for $M$ to be a causal mechanism. Thus, it is no surprise that this simple ad hoc approach is one of the dominant ways to estimate causal mechanisms, accounting for roughly 40% of the papers in our literature review.
@@
In this paper, we show that, in an experimental context, intermediate outcome tests under randomization of treatment alone can neither establish nor rule out the existence of an indirect effect of treatment on the outcome through that mediator. Papers that imply otherwise engage (often unknowingly) in what we call assumption smuggling—presenting suspect conclusions that rely on strong hidden assumptions alongside well-grounded findings that depend on the more reasonable, explicitly stated assumptions. We establish the IOT’s inability to inform mechanisms using sharp nonparametric bounds that contain all values of the indirect effect consistent with the observed data and the maintained assumptions. Under randomization of treatment alone, these bounds are never informative about the sign of the indirect effect for two reasons. First, the average effect of the treatment on the mediator might mask heterogeneous treatment effects that are important to the causal mechanism. Second, establishing indirect effects requires strong evidence or assumptions about the effect of the mediator on the outcome at the individual level.
@@
To help researchers navigate this, we further introduce an assumption that, though very rarely discussed in practice, does enable the IOT procedure to have a meaningful interpretation. Specifically, we investigate a monotonic mediator response assumption wherein the units’ value of the mediator can only be affected by treatment in one direction. Under this assumption, we show that a lack of treatment effect on the mediator can, in fact, rule out indirect effects through that mediator. We also derive the nonparametric sharp bounds for the average indirect effect under this assumption, showing that IOTs can “approximately” rule out mechanisms in that small (zero) effects of the treatment on the outcome imply small (zero) indirect effects. Of course, monotonicity is a strong, untestable assumption that will frequently fail to hold in applied practice, and there is no design-based way to ensure it holds. Thus, it is critical for researchers to state this assumption clearly and to interrogate it if they wish to use IOTs to assess causal mechanisms at all. We present one falsification test for this assumption that relies on nonparametric tests of effect heterogeneity, such as causal forests.
@@
Can IOTs ever establish the sign of an indirect effect? In some substantive settings, researchers may care more about the directional relations between the treatment and the mediator, rather than pinpointing the exact magnitude of the effect. We also show that, unfortunately, it is nearly impossible to establish the sign of an indirect effect even under very strong assumptions. In particular, we show that even if we assume a positive average effect of the mediator on the outcome, the sharp nonparametric bounds on the indirect effect will never be informative about the sign of that effect. Thus, while the monotonic mediator response assumption may allow researchers to rule out potential pathways, no intermediate outcome test can establish the simple existence of a causal mechanism even if we make heroic assumptions about the mediator-outcome relationship.
@@
The contributions of this paper are twofold. First, we highlight for applied researchers that this ad hoc, widely used approach to analyze causal mechanisms actually relies on strong and usually unstated assumptions; relaxing the assumptions, as we show, can render substantive conclusions invalid. Second, by clarifying the assumptions needed, we provide guidance for interpreting these IOTs in terms of the substantive application. The takeaway message of this paper reinforces what has been argued elsewhere: establishing causal mechanisms is difficult and requires strong assumptions regardless of the approach. Researchers may not avoid this difficulty even when using IOTs.
@@
Our work contributes to a broader literature on causal mechanisms and (in)direct effects in multiple fields. Many scholars have formalized the concept of direct and indirect effects in terms of potential outcomes and have derived assumptions that can point identify these quantities (Pearl, 2001; Imai, Keele and Yamamoto, 2010). Most similar to our own work is Imai, Tingley and Yamamoto (2013), which derived nonparametric bounds for indirect effects under several different experimental designs, and Glynn (2012), which showed how individual-level heterogeneity in treatment effects renders many indirect effect tests ineffective. We expand on their results by showing how different causal assumptions can change the bounds for a single experiment. Several studies have proposed alternative ways to assess causal mechanisms, such as implicit mediation (Bullock and Green, 2021) and effect heterogeneity (Fu and Slough, 2024). However, these papers show that these alternatives also rely on strong assumptions. An alternative definition of causal mechanisms centers on a quantity known as the “controlled direct effect” (CDE), which captures the average direct effect of a treatment variable while holding the mediator fixed at a given value (Goetgeluk, Vansteelandt and Goetghebeur, 2008; Acharya, Blackwell and Sen, 2016, 2018; Zhou and Wodtke, 2019; Blackwell and Strezhnev, 2022). While identifying the CDE requires weaker assumptions than mediation, many mediation-skeptical scholars object to these approaches because they require us to identify the effect of the mediator on the outcome. [^1] Some recent work investigates how to handle multiple mediating variables (Imai and Yamamoto, 2013; VanderWeele, Vansteelandt and Robins, 2014; Daniel et al., 2015; Zhou and Yamamoto, 2023). These approaches typically rely on stronger identification assumptions than the single-mediator case and are thus less common in political science. For these reasons, we focus on the single-mediator setting.
Two recent papers are directly related to our setting. In an applied setting of human rights treaties, Strezhnev, Kelley and Simmons (2021) explores a sensitivity analysis procedure for estimating indirect effects under a set of linear models, which is a familiar setting for many empirical scholars. Our results, on the other hand, are fully nonparametric and do not depend on functional form assumptions. Kwon and Roth (2024) proposes a way to test whether a given mediator explains all of the effect of treatment, which they call the “sharp null hypothesis of full mediation,” and also highlight the importance of a monotonicity assumption on the effect of treatment on the mediator. Their approach focuses on a novel quantity of interest (the sharp null), whereas we show what researchers can learn about more traditional causal mechanism estimands.
@@
This paper proceeds as follows. In Section 2, we discuss the setting of our two empirical applications, showing how researchers use IOTs in practice. Since the phrase “causal mechanisms” means different things to different scholars, we next lay out our preferred notation and definitions for causal effects and causal mechanisms in Section 3. In Section 4, we formally analyze intermediate outcomes tests using a principal stratification approach and derive sharp nonparametric bounds under different assumptions. Section 5 applies these bounds to two recent studies, showing the limited information these tests provide about indirect effects. Section 6 discusses how scholars might consider mechanisms when they cannot meet the assumptions we discuss in this paper. Section 7 concludes and describes avenues for future research.
2 Two Motivating Applications
@@ 2.1 Reducing Outgroup Prejudice
An important literature in political psychology examines intergroup prejudices and evaluates how different individual-oriented interventions might durably reduce such prejudices. An approach that has garnered empirical support in a growing literature is the use of interpersonal conversations (Galinsky and Moskowitz, 2000; Bruneau and Saxe, 2012; Broockman and Kalla, 2016; Adida, Lo and Platas, 2018; Simonovits, Kezdi and Kardos, 2018; Audette, Horowitz and Michelitch, 2020; Lowe, 2021; Williamson et al., 2021), but an unresolved question in this recent work is which “narrative strategies” are most effective at reducing prejudice. For example, an interpersonal conversation that emphasizes taking the perspective of an outgroup member (“perspective taking”) may yield different outcomes from one that simply describes the negative experiences of the outgroup (“ perspective getting”).
@@ 2.2 Transitional Justice Museums and Support for Democracy
Many scholars have credited transitional justice policies with facilitating democratic consolidation, reconciliation, and peace-building in post-conflict societies (De Brito, Enríquez and Aguilar, 2001; Horne, 2014; Nalepa, 2010). Less is known, however, about how such policies operate at the individual level and through what mechanisms. In a field experiment in Chile, Balcells, Palanza and Voytas (2022) study one particular instance of transitional justice policy: museums. Researchers randomized students into either a control condition or a treatment condition consisting of an hour-long tour of a museum memorializing victims of General Augusto Pinochet’s dictatorship. The authors then fielded a survey about beliefs about democracy and found that treated students “display greater support for democratic institutions, are more likely to reject institutions associated with the repressive period, and are more supportive of restorative transitional justice policies" (Balcells, Palanza and Voytas, 2022, p. 496).
@@
Lastly, while we do not evaluate all mediators and outcomes proposed in the original studies given their richness, we highlight a selected set of variables in line with the main theoretical arguments of the original authors. Table 1 lists the mediators and outcomes in each of the examples discussed in this section that we will be focusing on in the application section. Further discussions on why we select these variables can be found in Section 5.
Example Treatment Mediator Outcome
Kalla and Broockman (2023) Perspective-getting (i.e., reading a story about an outgroup member and then summarizing the story) during canvassing Attributional thinking; Entativity; Reactive empathy; Self-outgroup merging Reduced prejudice against transgender people and undocumented immigrants
Balcells, Palanza and Voytas (2022) Museum tour about victims of Pinochet’s dictatorship Negative emotions elicited during the tour (e.g., feeling tense, scared, guilty, etc.) Increased support for democratic institutions in Chile
Table 1: Summary of Treatments, Mediators, and Outcomes in the Motivating Applications
@@ 3 Causal Mechanisms: Notation and Quantities of Interest
With this notation, it is possible to define potential outcomes that involve “natural” values of the mediator, such as $Y_{i}(1,M_{i}(0))$, which is the outcome that unit $i$ would take if it was assigned to treatment but had received its mediator value under the control condition. These types of potential outcomes are important to mediation, but they create challenges in estimation and interpretation due to their “cross-world” nature. In particular, a unit cannot simultaneously be subject to treatment and control conditions at the same time, so $Y_{i}(1,M_{i}(0))$ is not observable even in principle.
@@ 3.1 Identifying assumptions
Random treatment assignment in the experimental setting can justify Assumption 1. This assumption is more difficult to justify in observational studies, but a large body of work focuses on so-called “natural” experiments that could make this assumption believable. Under this assumption, we can identify the ATE as the average of covariate-specific effects,
@@
Many authors have expressed skepticism about Assumption 2 since it requires that there are no (measured or unmeasured) post-treatment confounders for the relationship between the mediator and the outcome. [^2] One alternative approach that has garnered some attention in recent years involves estimating a quantity known as the controlled direct effect (CDE), which captures the causal effect of a treatment when the mediator is fixed to a given value. This quantity is attractive because it is identified under a weaker assumption than Assumption 2, one that allows for the existence of observed “intermediate confounders” (confounders that are affected by the treatment and which affect the mediator and outcome). A drawback of this approach is that it does not quantify the extent of mediation through $M_{i}$ and only enables a researcher to rule out the existence of alternative mediators other than $M_{i}$ (see Acharya, Blackwell and Sen (2016)). This condition is often difficult to sustain when the mediator is observed at its natural value after treatment, as is commonly the case. Indeed, while many are willing to believe the identifying assumptions for treatment, assuming the same condition for the mediator is much less plausible.